Suppose that out of 1200 students at your university, 100 are on a break and not taking any classes this semester. Which of the following mathematical expressions describes the standard deviation of the sampling distribution of the proportion of 50 students who are on a break?

Select one:

a. √100×120050

b. √10050×1200

d. √(100/1200)×(1100/1200)/50
Answer #1

Answer

we know that the formula for the standard deviation of proportion is given as

sigma = sqrt{(p*(1-p))/n}

where p = 100/1200 and n = 50

setting the given values, we get

σ-V( (100/12001 (100/1200))))/50

we can write 1 _ (100/1 200 ) (1200-100)/1200-1100/1200

σ = V(( 10011200) * (1 100/1200))/50

So, option D is correct answer